OpenStudy (anonymous):

Factor the expression y^2+18y+81

sam (.sam.):

First think of "something" multiply "sonething" =81 and also adds up to 18

OpenStudy (anonymous):

9x9=81 9+9=18

sam (.sam.):

yes,so, (y+9)(y+9)

OpenStudy (anonymous):

Got it thanks

sam (.sam.):

yw

OpenStudy (callisto):

since 81 = 9^2 and the middle term is 2(9)(y) = 18y You can use the identity: (a+b)^2 = a^2 + 2ab + b^2 where in this case, a =y , b=9